OpenStudy (anonymous):

find the absolute value of the difference of the solutions to the equation 3^2x -10(3^x) +9=0.

OpenStudy (turingtest):

let u=3^x

OpenStudy (anonymous):

@TuringTest then what would the equation be? and what would you do next?

OpenStudy (turingtest):

if u=3^x what is u^2=?

OpenStudy (anonymous):

@TuringTest would it be 3^2x?

OpenStudy (turingtest):

yep :) so now make the substitutions into the equation and you will get a regular quadratic

OpenStudy (anonymous):

@TuringTest the the equation would be U^2-10U+9 and then you factor? i have no clue!

OpenStudy (turingtest):

what two numbers multiply to make 9 and add to make -10 ?

OpenStudy (anonymous):

@TuringTest i know how to factor and i got (u-9)(u-1) then what

OpenStudy (turingtest):

the zero factor property states that\[(u-9)(u-1)=0\]implies that either\[u-9=0~~~~or~~~~u-1=0\]or both solve those two equations separately, then sub back in for u

OpenStudy (anonymous):

@TuringTest can you show me how to do that im confused

OpenStudy (turingtest):

can you solve u-9=0 for u ? I sure hope so

OpenStudy (anonymous):

yes so you put those answers back in to the equation for X then what @TuringTest

OpenStudy (turingtest):

what do you mean "put them in the equation for x" ? We are solving for u, not x yet what are the two solutions for u ?

OpenStudy (anonymous):

9 and 1 @TuringTest

OpenStudy (turingtest):

yes, so we have u=1 and u=9 now sub back in our original expression for u and what do we get?

OpenStudy (anonymous):

would the equations then be 3^18 -10(3^9)+9 and 3^2 -10(3) +9 @TuringTest

OpenStudy (turingtest):

you plugged in the values for u into where the x goes, that's not right. all I'm asking you to do is put the expressions u=1 and u=9 back in terms of x

OpenStudy (turingtest):

you also don't need to write @turingtest every time...

OpenStudy (turingtest):

they want to absolute value of the difference of the two solutions we have the two solutions in terms of u undo our original sub for u to find what x is supposed to be

OpenStudy (anonymous):

9=3^x 1=3^x?

OpenStudy (turingtest):

yes, excellent can you solve those for x?

OpenStudy (anonymous):

x=2,0

OpenStudy (turingtest):

great, and the absolute value of the difference of those two solutions is...?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

which is your answer, nice job :)

OpenStudy (anonymous):

thanks so much!:)

OpenStudy (turingtest):

very welcome!